40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 7 11:25:31 2020, gautam, Update, SUS, MC1 suspension glitchy again 
    Reply  Wed Dec 9 16:22:57 2020, gautam, Update, SUS, Yet another round of Sat. Box. switcharoo WFS2.pngWFS_lineNoise.pngWFSchar.pdf
       Reply  Sun Jan 3 16:26:06 2021, Koji, Update, SUS, IMC WFS check (Yet another round of Sat. Box. switcharoo) Screen_Shot_2021-01-03_at_17.14.57.png
    Reply  Thu May 13 11:55:04 2021, Anchal, Paco, Update, SUS, MC1 suspension misbehaving Screenshot_from_2021-05-13_09-50-24.pngMC1_Glitches_Invest2.pdf
       Reply  Thu May 13 19:38:54 2021, Anchal, Update, SUS, MC1 Satellite Amplifier Debugged MC1_UL_Channel_Fixed.png
          Reply  Mon May 24 19:14:15 2021, Anchal, Paco, Summary, SUS, MC1 Free Swing Test set to trigger 
             Reply  Tue May 25 10:22:16 2021, Anchal, Paco, Summary, SUS, MC1 new input matrix calculated and uploaded SUS_Input_Matrix_Diagonalization.pdf
          Reply  Thu Jun 17 11:45:42 2021, Anchal, Paco, Update, SUS, MC1 Gave trouble again SummaryScreenShot.pngMC1_LL_SENSOR_DEAD.png
             Reply  Thu Jun 17 16:37:23 2021, Anchal, Paco, Update, SUS, c1susaux computer rebooted 
                Reply  Tue Jun 22 11:56:16 2021, Anchal, Paco, Update, SUS, ADC/Slow channels issues CDS_FE_Status.png
                   Reply  Tue Jun 22 16:52:28 2021, Paco, Update, SUS, ADC/Slow channels issues shake_and_damp.png
                      Reply  Wed Jun 23 09:05:02 2021, Anchal, Update, SUS, MC lock acquired back again 
                         Reply  Thu Jun 24 16:40:37 2021, Koji, Update, SUS, MC lock acquired back again P_20210624_163641_1.jpg
Message ID: 16223     Entry time: Thu Jun 24 16:40:37 2021     In reply to: 16222
Author: Koji 
Type: Update 
Category: SUS 
Subject: MC lock acquired back again 

[Koji, Anchal]

The issue of the PD output was that the PD whitened outputs of the sat amp (D080276) are differential, while the successive circuit (D000210 PD whitening unit) has the single-ended inputs. This means that the neg outputs (D080276 U2) have always been shorted to GND with no output R. This forced AD8672 to work hard at the output current limit. Maybe there was a heat problem due to this current saturation as Anchal reported that the unit came back sane after some power-cycling or opening the lid. But the heat issue and the forced differential voltage to the input stage of the chip eventually cause it to fail, I believe.

Anchal came up with the brilliant idea to bypass this issue. The sat amp box has the PD mon channels which are single-ended. We simply shifted the output cables to the mon connectors. The MC1 sus was nicely damped and the IMC was locked as usual. Anchal will keep checking if the circuit will keep working for a few days.

Attachment 1: P_20210624_163641_1.jpg  1.497 MB  | Hide | Hide all
P_20210624_163641_1.jpg
ELOG V3.1.3-